In the temperature range between 0


0C and 7000C the resistance R [in ohms] of a certain platinum resistance


thermometer is given by


R = 10 + 0.04124T − 1.779 × 10−5T

2

where T is the temperature in degrees Celsius. Where in the interval from 0


0C to 7000C is the resistance of the

thermometer most sensitive and least sensitive to temperature changes? [Hint: Consider the size of dR/dT in the

interval 0 ≤ T ≤ 700.].


1
Expert's answer
2021-10-20T15:26:49-0400

The derivative of the function measures the rate at which function changes with respect to the change of the variable. So, to find where the given function is more/less sensitive to the changes of time we should find maximum/minimum of its derivation.

"R = 10 + 0.04124T \u2212 1.779 \u00d7 10^{\u22125T^2}"

"{\frac {dR} {dT}}=0.04124+17.79*ln(10)*T*10^{-5t^{2}}"

Now to find extremums of this function we must find it's derivative

"{\frac {d^{2}R} {dT^{2}}}=17.19*ln(10)*10^{-5t^{2}}-177.9*ln^{2}(10)*t^{2}*10^{-5t^{2}}"

Next step is to find the critical points:

"17.19*ln(10)*10^{-5t^{2}}-177.9*ln^{2}(10)*t^{2}*10^{-5t^{2}}=0"

"17.19*ln(10)*10^{-5t^{2}}(1-10*ln(10)*t^{2})=0"

"10^{-5t^{2}}(1-10*ln(10)*t^{2})=0"

Since "10^{-5t^{2}}"is never equal to 0, then

"1-10*ln(10)*t^{2}=0\to t = {\frac 1 {\sqrt{10*ln(10)}}}"

We have found the critical points, now we should verify the max and min at t є [0;700]

Let "{\frac {dR} {dT}}=f(t)"

"f(0) = 0.04124"

"f({\frac 1 {\sqrt{10*ln(10)}}})= 5.191"

"f(700) = 0.04124 + 17.79+ln(10)*700*10^{-5*700^{2}}"

"7.79+ln(10)*700*10^{-5*700^{2}}" is a bit higher than 0, than

"f(0) < f(700) < f({\frac 1 {\sqrt{10*ln(10)}}})"

The most sensitive thermometer is at point "T = {\frac 1 {\sqrt{10*ln(10)}}}" , less sensitive is at point T = 0